OpenStudy (anonymous):

If each score in a set of data scores were increased by 8 points, how would this affect the mean and the standard deviation?

OpenStudy (amistre64):

im gonna predict it moves the mean by 8 as well 2+3+3+7+10 = = 25/5 = 5 10+11+11+15+18 = 65/5 = 13 = 5+8

OpenStudy (amistre64):

the standard deviation measures the spread of the data; and the spread remains the same .... its just moved to the right by eight